OpenStudy (anonymous):

Find the sum of the first 30 terms of the sequence below. an=3n+2

OpenStudy (midhun.madhu1987):

When n=1, an = 3*1 + 2 = 5 when n=2, an = 3*2 + 2 = 8 when n = 3, an = 3*3 + 2 = 11

OpenStudy (midhun.madhu1987):

So the arithmatic sequence is 5,8,11,14......... Sum to n terms of an arithmatic sequence is given by: